OpenStudy (anonymous):

I was working on an MVC problem but forgot my trig: \(\sin(\pi-x) = ?\). It's like negative cosine or something, right?

OpenStudy (anonymous):

I think it's regular sine....

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yeah probably b/c sin(pi-x) = -sin(x-pi), and that's just flip n' shift

OpenStudy (anonymous):

i means yes, it is sine

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

@satellite73 while you're here - is there anyway of simplifying \[\sin(\arcsin(c_1\sin\Phi)+\alpha)\] Maybe with the sum of angles for sines we can cancel out a sin/arcsin term - or will we be stuck with a cos(arcsin) term?

OpenStudy (anonymous):

nvm ill ask a question seperately.